OpenStudy (anonymous):

what is 2/5 + 1/8 and what is 4/8 + 1/2

OpenStudy (anonymous):

2/5 + 1/8 = 21/40 4/8 + 1/2 = 1

OpenStudy (anonymous):

thanks

OpenStudy (unklerhaukus):

so to find 2/5 + 1/8 we first need common demonstrator (a number that is a dividable by 5 and 8 , the obvious candidate is just 5 x 8 which is 40 2/5 = 16/40 and 1/8 =5/40 now we have a common denominator we can simply add the numerators...

OpenStudy (anonymous):

It's much easier to just multiply the denominators and then cross multiply, then add the top numbers of the fraction for your final answer. Example: 2/5 + 1/8 5x8 = 40 5x1=5 2x8=16 16+5=21 Final result: 21/40